Suppose you bought 550 shares of stock at an initial price of $51 per share. The stock paid a dividend of $0.54 per share during the following year, and the share price at the end of the year was $52.

What is the capital gains yield and dividend yield?
Answer #1

Answer - Capital gain yield = 1.96% , Dividend yield = 1.06%

Reason -

Capital gain yield = (year end price - initial price) ÷ initial price * 100

= (52-51) ÷ 51 * 100

= 1 ÷ 51 * 100

= 1.960784313 or 1.96% (rounded off)

Dividend yield = dividend per share ÷ initial price * 100

= 0.54 ÷ 51 * 100

= 1.058823529 or 1.06% (rounded off)

If the solution helped, please give it a thumbs up. Thank you.